how do i find the appropriate formula to find the surface area of the room 9ft x 12ft?

Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(69443)

Area = length times width
= 9*12
= 108 square feet

Answer by rfer(16322)

A=L*W
A=9*12
A=108 sq ft